OpenStudy (anonymous):

Simplify. 20. (n^2 – 3n + 1) – (4n^2 + n – 5) 21.4y^2z(2y + 3z – 5) 22.(x – 8)(x + 5) Please help, I need this by the end of the day, or may not pass. Please :(

OpenStudy (pfenn1):

20.\[(n^2 – 3n + 1) – (4n^2 + n – 5)=n^2-3n+1-4n^2-n+5\]Combine like terms.

OpenStudy (anonymous):

I need the answers please, I dont understand and this is my last project of the year. If you could do that I would thanks you times 10.

OpenStudy (pfenn1):

\[(1-4)n^2+(-3-1)n+(1+5)=?\]

OpenStudy (anonymous):

I dont know, that's why i posted it on here. Please sir, I need this fast.